OpenStudy (anonymous):

What is the real roots of x^3-1=0

hero (hero):

x^3-1= 0 x^3 = 1 x^3 = 1^3 cube root(x^3) = cuberoot(1^3) x = 1

OpenStudy (anonymous):

x = 1

OpenStudy (dumbcow):

x^3 - 1 = 0 (x-1)(x^2+x+1) = 0 x-1 = 0 --> x = 1 x^2 + x +1 = 0 --> x = (-1+-isqrt3)/2

hero (hero):

dumb cow, it only wanted the real roots